# A Comprehensive Survey in LLM(-Agent) Full Stack Safety: Data, Training and Deployment

### v1

@@ 4.4.2 Provably Safe AI System
Provably safe AI systems represent an emerging paradigm that aims to ensure that advanced AI operates within rigorous, formally verifiable safety bounds. Some researchers argue that only by embedding mathematically verified safety proofs into AI architectures can we guarantee that such systems will never deviate into harmful behaviors [347]. This formal approach contrasts sharply with traditional empirical testing and red-teaming methods, which often fail to uncover all failure modes in complex or adversarial environments. The achievement of provable safety requires the integration of several key components [348]:
@@ 5 Safety in Model Editing & Unlearning
Concretely, model editing [370, alphaedit] and unlearning [371, 372, 373, 374] can be understood as methods to efficiently modify model parameters during deployment to enhance the model’s security and privacy. To better reflect the comprehensiveness of our survey, we have included relevant literature on the safety of editing (Section 5.1) and unlearning (Section 5.2). It is noteworthy that there exists a certain degree of technical overlap between model editing and unlearning. To provide a clearer and more precise exposition, we focus model editing on addressing knowledge conflicts within the model, while unlearning is primarily concerned with the erasure of knowledge to ensure privacy protection.
@@ 5.1 Safety in Model Editing
Gradient. Early approaches [376, 377, 383] advocate that the updating of knowledge in the LLMs is accomplished by modifying the gradient of the LLM. However, since gradient-based methods are too complex and suffer from pattern collapse, it is gradually being replaced by other research lines [384, 385].
@@
Locate-then-edit. Locate-then-edit methods, represented by RoME[382], MEMIT[386] and AlphaEdit[387], localizing knowledge storage-related neurons by causal tracing, achieving knowledge editing by modifying these neurons, have made breakthroughs in recent years [388, 389, 390]. The locate-then-edit approach has been proven to be effective in updating specific factual knowledge in the LLM [387]. Thus it is widely used to edit the security of LLMs [391, 392]. In the following section, we will focus on the application of the locate-then-edit approach to the security domain.
@@
Model editing methods have made big strides in red team, making them an effective means of injecting risk content into safely aligned models. We summarize the mainstream editing for attacks and defenses in Table V. Against model editing attacks, no research has been done to make a specific defense against such attacks, so further exploration in this area is an important research topic.
@@ 5.2 Safety in Unlearning
LLMs have demonstrated remarkable capabilities in various tasks, but their training on vast and often unfiltered datasets from the Internet inevitably leads to the absorption of unsafe information [399, 400, 401, 402, 403, 404]. This includes biases [405], stereotypes [406], toxic language [407], misinformation [408, 409, 410], and even private data [69]. Therefore, LLM unlearning is crucial for ensuring their safe and responsible deployment [411, 372], as shown in Figure 6. Unlearning, in this context, refers to the process of selectively removing or mitigating the influence of specific knowledge, behaviors, or data points from a trained LLM [412, 413, 414, 415, 416, 417, 418, 419, 420, 421]. Two primary paradigms have emerged to achieve this: parameter-adjusting methods, which modify the model’s internal weights, and parameter-preserving methods, which intervene externally without altering the core model architecture (refer to Figure 6).
Parameter-Adjusting Unlearning. The first paradigm, which involves adjusting the model’s parameters, is characterized by its direct intervention in the model’s internal structure. This approach typically requires retraining or fine-tuning the model on a curated dataset designed to counteract the unsafe knowledge or behavior that needs to be unlearned. Techniques such as Gradient Ascent [422] and its variations [423] are commonly employed. While traditional fine-tuning using cross-entropy loss is prevalent, more specialized loss functions have been proposed to enhance the control over the outputs of unlearned models, such as KL minimization [424, 425, 426] and the IDK loss function [427]. Additionally, recent work [428] has reframed LLM unlearning as a preference optimization problem [429], utilizing Negative Preference Optimization loss to improve the unlearning process. Given the recent powerful multimodal perception and generation nature of LLMs, MMUnlearner [430] proposes to reformulate the setting of multimodal unlearning, which aims at erasing the unwanted visual concept but still preserving textual knowledge. Based on existing multimodal LLM-based unlearning benchmarks [431, 432, 433], SafeEraser [434] further incorporates unlearning mechanism and evaluation into multimodal LLM safety, via introducing Prompt Decouple Loss and a new metric called Safe Answer Refusal Rate.
@@ 5.3.1 Model Editing
In the era of sophisticated safety alignment, model editing addresses a critical niche. While safety fine-tuning establishes systematic safeguards through periodic retraining, it struggles with emergent, context-sensitive risks (e.g., geopolitical shifts or cultural updates) that evolve faster than retraining cycles. As LLMs scale, the intervals between alignment updates widen, creating safety gaps exacerbated by catastrophic forgetting risks. Model editing bridges these gaps through rapid surgical interventions — executing updates orders of magnitude faster than alignment procedures — by modifying specific unsafe knowledge or concepts, all while preserving general model stability. In summary, while safety fine-tuning remains essential for systematic alignment, model editing addresses three fundamental limitations in the current era:
@@
  • More Hidden Backdoor: By precisely modifying targeted parameters without perturbing unrelated knowledge, edited backdoors evade traditional detection methods that monitor broader model behavior;
@@ 6.1.1 Attack in Deployment
Model Extraction Attacks. Model extraction attacks aim to steal a deployed language model, which only provides an Application Programming Interface (API) that processes text input (i.e., a prompt) and returns generated outputs. Carlini et al. [462] conducted the first model-stealing attack against a black-box large language model by targeting its embedding projection layer. Building on this, Finlayson et al. [463] further investigated the risk of stealing embedding dimensions by exploiting the softmax bottleneck. Another line of research explores model extraction in a gray-box setting. For instance, Zanella et al. [464] demonstrated the feasibility of stealing high-fidelity language models when given access to a frozen or fine-tuned encoder.
@@
  • Methods. In the beginning work, [468] proposes the first MIA with MIN-KPROB, which identifies examples that contain few outlier words with low probabilities as non-members. Afterward, [469] proposes MIN-K%++, which simulates the membership inference into identifying local maxima. Some works reveal that the success of MIAs against LLMs may be due to sampling non-members from different distributions. Thus, [470] proposes Blind attack, which conducts MIA by applying a threshold and completely ignores the target model. [471] selectively combines the existing MIAs and aggregates their scores to perform a statistical test. [472] identifies the membership of a verbatim text by constructing paraphrased options (with another proxy model) and asking the target LLM for true verbatim. [473] examines the relative change in conditional log-likelihoods when prefixing target data points with non-member context. [474] proposes to generate noisy neighbors for a target sample by adding stochastic noise in the embedding space. [475] train a neural network to capture variations in output probability distributions between members and non-members.
@@
  • Factor Impact. Duan et al. [110] find that the existing MIAs work poorly on LLM due to massive training data and near-one-epoch training. Li et al. [487] clarify the impact of fine-tuning and evaluation metrics and propose a three-phase framework (i.e. training, simulation, and confidence calculation) to assess membership leakage. Kandpal et al. [86] find that duplication of training data highly extends the risk of MIA. Naseh et al. [488] validate that using synthetic data in membership evaluations may lead to false classification results.
@@
AntoDAN-Trubo [498] Yes 2024 LLM-based Black-box Harmbench Llama-2, Gemma, GPT-4, Gemini ASR, StrongREJECT
@@
Jailbreak Attacks. Jailbreak attacks aim to induce the large language model to generate unsafe content like violence [241]. Jailbreak attacks focus on bypassing the safety rules, including system safety prompts and safety filters, while prompt injection attacks target all system prompts. Lots of literature has studied the vulnerability of LLM, where different terms, including “jailbreak attack” and “red-teaming”, all point to the same safety vulnerability that generates unsafe content. We classify them into two main categories, i.e. optimization-based and strategy-based.
Strategy-based jailbreaks figure out novel strategies or templates to generate one adversarial prompt at a heat to test LLMs’ vulnerabilities, which are pre-defined. Thus, the generated prompt is non-evolvable. Specifically, useful strategies include persuasion [516], role-playing [517, 518, 519, 520], cipher [521, 522], ASCII [523], long-context [524], low-resource language [525], in-context malicious demonstration [526], overloaded logical thinking [527], misspelling [528], multi-language mixture [529], rephrasing [530, 497, 531, 532], competing objectives and generalization mismatch [533], zero-shot generation [534], personal modulation [535].
@@
  • LLM-based Optimization. PAIR [242] constructs a system prompt and uses an Attacker LLM to generate and revise adversarial prompts. It also uses a judge model to assess the feedback from the victim, which is further utilized for revising the adversarial prompt. AutoDAN-A [493] utilizes crossover strategies and LLM-based mutation to revise adversarial prompts into stealthy sentences. AntoDAN-Trubo [498] proposes to find useful strategies by prompting an LLM automatically. ToA (Tree of Attack) [495] iteratively uses an LLM to transform the unsafe prompt into two variations and keeps the prompt variation that achieves a higher score. Xiao et al. [540] adopt a similar pipeline with PAIR [242] and introduce malicious content concealing and memory reframing. Puzzler [541] proposes defensive and offensive measures to conduct an indirect jailbreak. GPTFUZZER [542] starts from human-written prompts, and uses templates and mutation to rewrite unsafe prompts. ECLIPSE [543] uses an LLM as a suffix generator and optimizer. PAL [496] proposes an online proxy model (which is used for adversarial prompt generation) training pipeline.
@@
  • Different Settings. Some works also explore the risk of data leakage in novel settings. Wang et al. [512] studies the probability of data extraction in fine-tuning settings and Bargav et al. [558, 559] extract the training data by comparing the output difference before and after the fine-tuning. Jiang et al. [560, 561, 562] propose to extract the private Retrieval-Augmented Generation (RAG) documents. Peng et al. [563] extract the private RAG documents by poisoning in the fine-tuning process. Nasr et al. [101] explore the potential risk of data extraction for the aligned production language models. Panda et al. [564] extract the fine-tuning secret data by poisoning the pertaining dataset. Lu et al. [565] propose to extract PII from an aligned model with model merging. Chen et al. [566] find that fine-tuning can recover the forgotten PIIs in pretraining data. Panchendrarajan et al. [567] propose to extract the whole private training data in the fine-tuning process. Rashid et al. [568] propose selective weight tampering to explore PPI leakage in Federated Language Models. Dentan et al. [569] extract data from layout-aware document understanding models like unimodal or bimodal models.
@@
  • Different Applications. Leveraging the abnormally high token probabilities, some works utilize the memorization of LLMs to extract the fingerprint or steganography [570]. Al-Kaswan et al. [571] explore memorization in large language models for code and find that code models memorize training data at a lower rate than natural language models. Nie et al. [572] utilize the token-level features derived from the identified characteristics to decode the PII. Lehman et al. reveal the risk of Electronic Health Records leakage of LLMs [573]. Diera et al. [574] conducted experiments to assess the PII leakage of fine-tuned BERT models and found that Differential Privacy (DP) has a negative effect when deployed in fine-tuning. Zhang et al. [575] propose data extraction attacks against text classification with transformers. Huang et al. [576] propose an evaluation tool, i.e. HCR, to assess the PPI leakage in Neural Code Completion Tools.
@@
  • Factor Assessment. Some work studies the factors of data extraction, including decoding schemes, model sizes, prefix lengths, partial sequence leakages, and token positions [577, 578]. Yash et al. [579] explore the effects of prompt sensitivity and access to multiple checkpoints to extraction attacks. Staab et al. [580] construct a dataset consisting of real Reddit profiles to extract personal attributes. Xu et al. [581] conduct experiments to evaluate the factors of different suffix generation methods and different membership inference attacks in extraction performance. Karamolegkou et al. [582] evaluate the effect of model structure, data type, probing strategies, and metrics.
@@
Prompt Stealing Attacks. Given that crafting effective prompts requires significant engineering effort and can be considered valuable intellectual property (IP), prompt-stealing attacks aim to compromise this IP by reconstructing prompts from generated responses [513, 514, 515]. These generation effects are often showcased to attract potential prompt buyers. Sha et al. [513] pioneered this approach by collecting a dataset and training classifiers to predict prompt parameters—such as whether the prompt is direct, role-based, or in-context. They then used a large language model (LLM) to reconstruct the prompt. Similarly, Zhang et al. [514] trained an LLM on output-prompt pairs to directly infer the original prompt, while Yang et al. [515] leveraged generation differences to refine surrogate prompts. However, recovering the original prompt solely from the output is challenging. Out of this, Zheng et al. [583] propose a timing-based side-channel method to infer the prompt during inference.

### v2

@@ 4.4.2 Provably Safe AI System
Provably safe AI systems represent an emerging paradigm that aims to ensure that advanced AI operates within rigorous, formally verifiable safety bounds. Some researchers argue that only by embedding mathematically verified safety proofs into AI architectures can we guarantee that such systems will never deviate into harmful behaviors [371]. This formal approach contrasts sharply with traditional empirical testing and red-teaming methods, which often fail to uncover all failure modes in complex or adversarial environments. The achievement of provable safety requires the integration of several key components [372] as follows:
@@ 5 Safety in Model Editing & Unlearning
Concretely, model editing [394, 395] and unlearning [396, 397, 398, 399, 400, 401] can be understood as methods to efficiently modify model parameters during deployment to enhance the model’s security and privacy. To better reflect the comprehensiveness of our survey, we have included relevant literature on the safety of editing (Section 5.1) and unlearning (Section 5.2). It is noteworthy that there exists a certain degree of technical overlap between model editing and unlearning. To provide a clearer and more precise exposition, we focus model editing on addressing knowledge conflicts within the model, while unlearning is primarily concerned with the erasure of knowledge to ensure privacy protection.
@@ 5.1 Safety in Model Editing
Gradient. Early approaches [403, 404, 410] advocate that the updating of knowledge in the LLMs is accomplished by modifying the gradient of the LLM. A more recent study [411] revisits gradient-based fine-tuning and demonstrates strong performance through constrained optimization techniques. However, since gradient-based methods are too complex and suffer from pattern collapse, it is gradually being replaced by other research lines [412, 413].
@@
Locate-then-edit. Locate-then-edit methods, represented by RoME [409], MEMIT [414] and AlphaEdit [395], localizing knowledge storage-related neurons by causal tracing, achieving knowledge editing by modifying these neurons, have made breakthroughs in recent years [415, 416, 417]. The locate-then-edit approach has been proven to be effective in updating specific factual knowledge in the LLM [395]. Thus it is widely used to edit the security of LLMs [418, 419]. In the following part, we will focus on the application of the locate-then-edit approach to the security domain.
@@
Model editing methods have made big strides in red team, making them an effective means of injecting risk content into safely aligned models. We summarize the mainstream editing for attacks and defenses in Table V and each row in the table represents distinct included content.. Against model editing attacks, no research has been done to make a specific defense against such attacks, so further exploration in this area is an important research topic.
@@ 5.2 Safety in Unlearning
LLMs have demonstrated remarkable capabilities in various tasks, but their training on vast and often unfiltered datasets from the Internet inevitably leads to the absorption of unsafe information [426, 427, 428, 429, 430, 431]. This includes biases [432], stereotypes [433], toxic language [434], misinformation [435, 436, 437], and even private data [71]. Therefore, LLM unlearning is crucial for ensuring their safe and responsible deployment [438, 399], as shown in Figure 6. Unlearning, in this context, refers to the process of selectively removing or mitigating the influence of specific knowledge, behaviors, or data points from a trained LLM [439, 440, 441, 442, 443, 444, 445, 446, 447, 448, 449]. Unlearning methods can be distinguished into two broad paradigms [450]: exact (certified) unlearning and heuristic (approximate) unlearning. Exact methods accurately identify poisoned data points or affected parameters, providing formal or statistical guarantees that the specified behaviors no longer influence the model. This typically requires certified retraining from scratch, removing the disallowed data entirely [451]. Two primary paradigms have emerged to achieve approximate unlearning: parameter-adjusting methods, which modify the model’s internal weights, and parameter-preserving methods, which intervene externally without altering the core model architecture (refer to Figure 6).
Parameter-Adjusting Unlearning. The first paradigm, which involves adjusting the model’s parameters, is characterized by its direct intervention in the model’s internal structure. This approach typically requires retraining or fine-tuning the model on a curated dataset, designed to counteract the unsafe knowledge or behavior that needs to be unlearned. It also encompasses methods that follow a locate-then-edit pipeline, where specific parameters associated with the target knowledge are identified and directly modified to achieve unlearning [449]. Techniques such as Gradient Ascent [452] and its variations [453] are commonly employed. While traditional fine-tuning using cross-entropy loss is prevalent, more specialized loss functions have been proposed to enhance the control over the outputs of unlearned models, such as KL minimization [454, 455, 456] and the IDK loss function [457]. Additionally, recent work [458] has reframed LLM unlearning as a preference optimization problem [459], utilizing Negative Preference Optimization loss to improve the unlearning process. In contrast to these training-intensive approaches, LaW [449] draws inspiration from model editing by identifying and removing knowledge associations embedded in MLP weights, aiming to eliminate targeted information with minimal impact on the model’s overall capabilities. Given the recent powerful multimodal perception and generation nature of LLMs, MMUnlearner [460] proposes to reformulate the setting of multimodal unlearning, which aims at erasing the unwanted visual concept but still preserving textual knowledge. Based on existing multimodal LLM-based unlearning benchmarks [461, 462, 463], SafeEraser [464] further incorporates unlearning mechanism and evaluation into multimodal LLM safety, via introducing Prompt Decouple Loss and a new metric called Safe Answer Refusal Rate.
@@
Although heuristic methods are far more scalable, their guarantees are only empirical. Closing this gap between certified safety and practical feasibility remains a central research challenge for the field.
@@ 5.3.1 Model Editing
In the era of sophisticated safety alignment, model editing addresses a critical niche. While safety fine-tuning establishes systematic safeguards through periodic retraining, it struggles with emergent, context-sensitive risks (e.g., geopolitical shifts or cultural updates) that evolve faster than retraining cycles. As LLMs scale, the intervals between alignment updates widen, creating safety gaps exacerbated by catastrophic forgetting risks. Model editing bridges these gaps through rapid surgical interventions — executing updates orders of magnitude faster than alignment procedures — by modifying specific unsafe knowledge or concepts, all while preserving general model stability. In summary, while safety fine-tuning remains essential for systematic alignment, model editing addresses four fundamental limitations in the current era:
@@
  • More Hidden Backdoor: By precisely modifying targeted parameters without perturbing unrelated knowledge, edited backdoors evade traditional detection methods that monitor broader model behavior.
@@ 6.1.1 Attack in Deployment
Model Extraction Attacks. Model extraction attacks aim to steal a deployed language model, which only provides an Application Programming Interface (API) that processes text input (i.e., a prompt) and returns generated outputs. He et al. and Peng et al. [492, 493, 494, 495] made a series of early efforts in launching model extraction or stealing attacks against LLMs (even deployed as a service) and proposed various defense mechanisms to mitigate such risks. Carlini et al. [496] conducted the model-stealing attack against a black-box large language model by targeting its embedding projection layer. Building on this, Finlayson et al. [497] further investigated the risk of stealing embedding dimensions by exploiting the softmax bottleneck. Another line of research explores model extraction in a gray-box setting. For instance, Zanella et al. [498] demonstrated the feasibility of stealing high-fidelity language models when given access to a frozen or fine-tuned encoder.
@@
  • Methods. [502] propose the first MIA with MIN-KPROB, which identifies examples that contain few outlier words with low probabilities as non-members. Afterward, [503] propose MIN-K%++, which simulates the membership inference into identifying local maxima. Some works reveal that the success of MIAs against LLMs may be due to sampling non-members from different distributions. Thus, [504] propose Blind attack, which conducts MIA by applying a threshold and completely ignores the target model. [505] selectively combine the existing MIAs and aggregate their scores to perform a statistical test. [506] identify the membership of a verbatim text by constructing paraphrased options (with another proxy model) and asking the target LLM for true verbatim. [507] examine the relative change in conditional log-likelihoods when prefixing target data points with non-member context. [508] propose to generate noisy neighbors for a target sample by adding stochastic noise in the embedding space. [509] train a neural network to capture variations in output probability distributions between members and non-members.
@@
  • Factor Impact. Duan et al. [117] find that the existing MIAs work poorly on LLM due to massive training data and near-one epoch training. Li et al. [521] clarify the impact of fine-tuning and evaluation metrics and propose a three-phase framework (i.e. training, simulation, and confidence calculation) to assess membership leakage. Kandpal et al. [87] find that duplication of training data highly extends the risk of MIA. Naseh et al. [522] validate that using synthetic data in membership evaluations may lead to false classification results.
@@
AutoDAN-Turbo [532] Yes 2024 LLM-based Black-box Harmbench Llama-2, Gemma, GPT-4, Gemini ASR, StrongREJECT
FlipAttack [533] Yes 2025 Rephrasing Black-box AdvBench, StrongREJECT GPTs, Claude 3.5 Sonnet, Llama 3.1 405B, Mixtral 8x22B ASR
Geneshift [534] Yes 2025 LLM-based Black-box AdvBench GPTs ASR
@@
Jailbreak Attacks. Jailbreak attacks aim to induce the large language model to generate unsafe content like violence [260]. Jailbreak attacks focus on bypassing the safety rules, including system safety prompts and safety filters, while prompt injection attacks target all system prompts. Lots of literature have studied the vulnerability of LLM, where different terms, including “jailbreak attack” and “red-teaming”, all point to the same safety vulnerability that generates unsafe content. We classify them into two main categories, i.e. optimization-based and strategy-based.
Strategy-based jailbreaks figure out novel strategies or templates to generate one adversarial prompt at a heat to test LLMs’ vulnerabilities, which are pre-defined. Thus, the generated prompt is non-evolvable. Specifically, useful strategies include persuasion [552], role-playing [553, 554, 555, 556], cipher [557, 558], ASCII [559], long-context [560], low-resource language [561, 562], in-context malicious demonstration [563], overloaded logical thinking [564], misspelling [565], multi-language mixture [566], rephrasing [567, 531, 568, 569], competing objectives and generalization mismatch [570], [wenjie: splitting sub-queries [571]], zero-shot generation [572], personal modulation [573].
@@
  • LLM-based Optimization. PAIR [261] constructs a system prompt and uses an attacker LLM to generate and revise adversarial prompts. It also uses a Judge model to assess the feedback from the victim, which is further utilized for revising the adversarial prompt. AutoDAN-A [527] utilizes crossover strategies and LLM-based mutation to revise adversarial prompts into stealthy sentences. AntoDAN-Trubo [532] AutoDAN-Turbo proposes to find useful strategies by prompting an LLM automatically. ToA (Tree of Attack) [529] iteratively uses an LLM to transform the unsafe prompt into two variations and keeps the prompt variation that achieves a higher score. Xiao et al. [578] adopt a similar pipeline with PAIR [261] and introduce malicious content concealing and memory reframing. Puzzler [579] proposes defensive and offensive measures to conduct an indirect jailbreak. GPTFUZZER [580] starts from human-written prompts, and uses templates and mutation to rewrite unsafe prompts. ECLIPSE [581] uses an LLM as a suffix generator and optimizer. PAL [530] proposes an online proxy model (which is used for adversarial prompt generation) training pipeline.
@@
  • Different Settings. Some works also explore the risk of data leakage in novel settings. Wang et al. [548] study the probability of data extraction in fine-tuning settings and Bargav et al. [596, 597] extract the training data by comparing the output difference before and after the fine-tuning. Jiang et al. [598, 599, 600] propose to extract the private Retrieval-Augmented Generation (RAG) documents. Peng et al. [601] extract the private RAG documents by poisoning in the fine-tuning process. Nasr et al. [107] explore the potential risk of data extraction for the aligned production language models. Panda et al. [602] extract the fine-tuning secret data by poisoning the pertaining dataset. Lu et al. [603] propose to extract PII from an aligned model with model merging. Chen et al. [604] find that fine-tuning can recover the forgotten PIIs in pretraining data. Panchendrarajan et al. [605] propose to extract the whole private training data in the fine-tuning process. Rashid et al. [606] propose selective weight tampering to explore PPI leakage in Federated Language Models. Dentan et al. [607] extract data from layout-aware document understanding models like unimodal or bimodal models.
@@
  • Different Applications. Leveraging the abnormally high token probabilities, some works utilize the memorization of LLMs to extract the fingerprint or steganography [608]. Al-Kaswan et al. [609] explore memorization in large language models for code and find that code models memorize training data at a lower rate than natural language models. Nie et al. [610] utilize the token-level features derived from the identified characteristics to decode the PII. Lehman et al. [611] reveal the risk of Electronic Health Records leakage of LLMs. Diera et al. [612] conduct experiments to assess the PII leakage of fine-tuned BERT models and found that Differential Privacy (DP) has a negative effect when deployed in fine-tuning. Zhang et al. [613] propose data extraction attacks against text classification with transformers. Huang et al. [614] propose an evaluation tool, i.e. HCR, to assess the PPI leakage in Neural Code Completion Tools.
@@
  • Factor Assessment. Some work studies the factors of data extraction including decoding schemes, model sizes, prefix lengths, partial sequence leakages, and token positions [615, 616]. Yash et al. [617] explore the effects of prompt sensitivity and access to multiple checkpoints to extraction attacks. Staab et al. [618] construct a dataset consisting of real Reddit profiles to extract personal attributes. Xu et al. [619] conduct experiments to evaluate the factors of different suffix generation methods and different membership inference attacks in extraction performance. Karamolegkou et al. [620] evaluate the effect of model structure, data type, probing strategies, and metrics.
@@
Prompt Stealing Attacks. Given that crafting effective prompts requires significant engineering effort and can be considered valuable intellectual property (IP), prompt-stealing attacks aim to compromise this IP by reconstructing prompts from generated responses [549, 550, 551]. These generation effects are often used to attract prospective prospective buyers. Sha et al. [549] pioneer this approach by collecting a dataset and training classifiers to predict prompt parameters—such as whether the prompt is direct, role-based, or in-context. They then used a large language model (LLM) to reconstruct the prompt. Similarly, Zhang et al. [550] trained an LLM on output-prompt pairs to directly infer the original prompt, while Yang et al. [551] leveraged generation differences to refine surrogate prompts. However, recovering the original prompt solely from the output is challenging. Out of this, Zheng et al. [621] propose a timing-based side-channel method to infer the prompt during inference.